OpenStudy (anonymous):

Find the equation of the line in the xy-plane that contains the point (-4, -5) and that is parallel to the line y + 2x = 3

OpenStudy (anonymous):

y=-2x-13?

OpenStudy (anonymous):

yep, that's what I got.

OpenStudy (anonymous):

Well, I can't draw this thing just gonna tell you how I got it.

OpenStudy (anonymous):

so y+2x=3 change it to slope-intercept form which becomes (y=mx+b) y=-2x+3 where m=slope and b=y-intercept since it's rise/run our slope rises by -2 and runs by 1 so that's how I graphed the first equation. for the second one the line that passes through (-4, -5) I just copied the slope of the first one until I reached the y-axis which hit -13 so y=-2x-13